40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10444     Entry time: Wed Sep 3 04:17:21 2014
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Y green ALS (not PDH) needs investigation 

Q put the X PDH box back, so that I could try locking, and remember which end is up after a week away.

I am unable to hold ALS comm/diff for any length of time. Only once today did I hold it through the FM3 boost turn-on.  So, I looked at the individual arms.

Xarm, even though it's the one that Q is seeing this saturation problem with, seems fine. 

Yarm however is having trouble holding lock for more than a few minutes at a time.  The green beam stays locked to the arm for ~infinity, so I'm not so worried about the PDH box right now.  If I look at the error and control points of the ALS digital servo, the Yarm is much more noisy above about 20 Hz.  Something that I might think of for this kind of mismatch at higher frequencies is poorly matched whitening / dewhitening, or none at all for the Yarm, however this doesn't look like that to me.  Based on the shape of the spectra, I don't think that we're running into ADC noise. For this plot, both arms are individually locked with ALS feeding back to the ETM, gain magnitude of 15 (Xarm gets a minus sign because of our temperature / beatnote moving direction convention), FMs 1,2,3,5,6 on.  Something that seems critical for getting the Yarm to have the FM3 boost without losing lock is having the SLOW temperature servos on for a little while so that the PZT output (as monitored on the temp servo screen) for the end lasers fluctuate around zero. Right now, both beatnotes are at about 62MHz, with an amplitude of about -31dBm.

Yarm_noisy_above_20Hz.pdf

I still need to do a somewhat more thorough investigation of what might be causing the Yarm locklosses.  Is the length-to-angle decoupling worse for ETMY than for ETMX?  Am I moving the arm length so far that the PZT can't follow within its actuation limits?  Does the Yend PDH box have a similar saturation to the Xend box, but somehow (a) worse, and (b) not as obvious so we didn't suspect it before? 

I need to put this plot into calibrated units, and also include the low frequency monitor that we have of the PDH error point (all of which are _DQ channels).

Things to do:

* Figure out Xend PDH box saturation issue.  Is Yend seeing same saturation in the variable gain amplifier?  We have 3 spares of these chips in the Plateau Tournant Bleu, if we need them. 

* Check Yarm ALS stability.  (NB:  The arms have been individually locked for the last 15 min or so while I've been writing, so maybe letting the slow servo settle is the key, and this is not something that needs work).

* Get CARM on DC Trans, DARM on AS55Q (after arm powers of about 1).  Can we see good REFL DC dip?  Should we try using just the transmission PD signal as the error signal for the CM board, if we aren't close enough to resonance to use REFL DC?

ELOG V3.1.3-